# Learning Interpretable Hierarchical Dynamical Systems Models from Time Series Data

### v1

@@ Abstract
In science, we are often interested in obtaining a generative model of the underlying system dynamics from observed time series. While powerful methods for dynamical systems reconstruction (DSR) exist when data come from a single domain, how to best integrate data from multiple dynamical regimes and leverage it for generalization is still an open question. This becomes particularly important when individual time series are short, and group-level information may help to fill in for gaps in single-domain data. At the same time, averaging is not an option in DSR, as it will wipe out crucial dynamical properties (e.g., limit cycles in one domain vs. chaos in another). Hence, a framework is needed that enables to efficiently harvest group-level (multi-domain) information while retaining all single-domain dynamical characteristics. Here we provide such a hierarchical approach and showcase it on popular DSR benchmarks, as well as on neuroscientific and medical time series. In addition to faithful reconstruction of all individual dynamical regimes, our unsupervised methodology discovers common low-dimensional feature spaces in which datasets with similar dynamics cluster. The features spanning these spaces were further dynamically highly interpretable, surprisingly in often linear relation to control parameters that govern the dynamics of the underlying system. Finally, we illustrate transfer learning and generalization to new parameter regimes.
@@ 1 Introduction
Foundation models for time series recently blossomed, based on the hope that generalizable temporal features could be extracted from multiple, diverse datasets (Yeh et al., 2023; Miller et al., 2024; Das et al., 2024; Ansari et al., 2024; Rasul et al., 2024). In scientific and medical applications, however, beyond mere time series forecasting, we are often interested in interpretable and mathematically tractable models that can be used to obtain mechanistic insights into the underlying system dynamics. Such models need to be generative in a dynamical systems sense (Durstewitz et al., 2023), i.e., when simulated, should produce dynamical behavior with the same long-term statistics as the observed system, a property that conventional time series models often lack (Gilpin, 2024). Moreover, accounting for the fact that different time series may be sampled from fundamentally different dynamical regimes, i.e. with different state space topology and attractor structure, may also help to improve generalization and transfer learning (Göring et al., 2024).
This is particularly important in areas like neuroscience (Zeidman et al., 2019) or mental health (Huys et al., 2016; Valton et al., 2020), where the data is not only expensive and difficult to collect (Koppe et al., 2021; Wen et al., 2017), hence comparatively sparse, but also subject to significant variability between individuals or systems (Meyer-Lindenberg, 2023; Fechtelpeter et al., 2024). The challenge in these settings is to develop models that can still identify and exploit commonalities across different datasets while being flexible enough to account for individual differences in dynamical regimes.
@@
Hierarchical models provide a natural solution by separating domain-general and domain-specific features. Such models facilitate the transfer of group-level information across individual datasets (Pan and Yang, 2010), avoiding overfitting and improving generalization to unobserved conditions with minimal data (Kirchmeyer et al., 2022). They further explicitly represent inter-domain variability, as often crucial in addressing scientific hypotheses, by unsupervised extraction of a common feature space within which individual datasets are embedded (Zeidman et al., 2019). While hierarchical modeling is a meanwhile established approach in statistics, including time series analysis (Berliner, 1996; Hyndman et al., 2011), it has so far hardly been applied to the more challenging problem of dynamical systems reconstruction (DSR) where we aim for generative models of the system dynamics that can be used for downstream analysis and simulation.
@@
In this work, we propose such a general framework for extracting hierarchical dynamical systems models from multi-domain time series data. Our approach combines low-dimensional, system- or subject-specific parameter vectors with high-dimensional, group-level weights to produce domain-specific recurrent neural networks (RNNs), which can be trained end-to-end using state-of-the-art techniques designed for DSR (Mikhaeil et al., 2022; Hess et al., 2023). Beyond the provision of a common feature space within which all observed systems are embedded, our approach infers system-specific DS models that serve as generative models for the latent processes underlying individual time series observations. We demonstrate the effectiveness of our method in several ways: By integrating data across systems, we enhance model accuracy on multiple short time series (transfer learning). The method further extracts highly interpretable low-dimensional feature spaces from the dynamical systems benchmarks and empirical datasets employed here, including a medical dataset of arterial pulse waves (intepretability). We show how - once a general DSR model has been learned - new models can be inferred that generalize to previously unobserved dynamical regimes with minimal new data (few-shot learning). Finally, we validate our framework on human electroencephalography (EEG) and functional magnetic resonance imaging (fMRI) data, and demonstrate how unsupervised classification in the learned low-dimensional spaces based on dynamical systems features profoundly outperforms all other unsupervised detection methods that harvest more ‘conventional’ time series features, providing a strong point for the DSR perspective on time series analysis.
@@ 2 Related Work
DSR is a rapidly growing field in scientific machine learning. DSR models may be considered a special class of time series (TS) models that, beyond mere TS prediction, aim to learn surrogate models of the data-generating process from TS observations. A proper DSR model needs to preserve the long-term temporal and geometrical properties of the original dynamical system (DS), i.e. its vector field topology and attractor structure, which then enables further scientific analysis (Brunton and Kutz, 2019; Durstewitz et al., 2023; Platt et al., 2023; Gilpin, 2024). A variety of DSR methods have been advanced in recent years, either based on predefined function libraries and symbolic regression, such as Sparse Identification of Nonlinear Dynamics (SINDy; Brunton et al. (2016); Kaiser et al. (2018)), based on Koopman operator theory (Azencot et al., 2020; Brunton et al., 2021; Naiman and Azencot, 2021), on universal approximators such as recurrent neural networks (RNNs; Gajamannage et al. (2023); Brenner et al. (2022); Hess et al. (2023)), reservoir computing (RC; Pathak et al. (2017); Platt et al. (2022; 2023)), or neural ordinary differential equations (Neural ODEs; Chen et al. (2018); Ko et al. (2023)), or on methods that sit somewhere in between universal and domain-specific, like physics-informed neural networks (PINNs; Raissi et al. (2019)). PINNs, like library-based methods, require sufficient domain knowledge to work well in practice (Fotiadis et al., 2023; Subramanian and Mahadevan, 2023; Mouli et al., 2023). To achieve proper reconstruction of a system’s long-term statistics and attractor geometry, often special, control-theoretic training techniques such as sparse or generalized teacher forcing (Brenner et al., 2022; Mikhaeil et al., 2022; Hess et al., 2023) or particular regularization terms that enforce invariant measures (Platt et al., 2023; Jiang et al., 2023; Schiff et al., 2024) are used. Models using these training techniques currently represent the state of the art in this field.
@@
Hierarchical models for representing multi-level dependencies or nested groups have a long history in statistics and machine learning (Laird and Ware, 1982; Goldstein, 1987; Gelman and Hill, 2006; McCulloch et al., 2011). The general idea of these approaches is to leverage shared group-level information while accounting for individual-level variation in an, ideally, interpretable way. Many such frameworks are Bayesian (Allenby et al., 2005; Gelman et al., 2013), informing individual-level distributions by group-level priors, and have been widely applied in psychology and cognitive science (Lee and Wagenmakers, 2013; Kruschke, 2015). In the time series domain, Zoeter and Heskes (2003); Bakker and Heskes (2007), for instance, introduced dynamic hierarchical models that combine individual time series with group-level dependencies. Hierarchical models have also been applied for forecasting chaotic time series (Matsumoto et al., 1998; Hyndman et al., 2011; Xu et al., 2020), or for extracting interpretable summary statistics from time series to capture inter-individual differences (Akintayo and Sarkar, 2018; Yingzhen and Mandt, 2018), for example for clustering dynamical patterns to identify neuron groups with similar response profiles (Lin et al., 2019). Much less work exists in the direction of hierarchical dynamical systems modeling. Roeder et al. (2019) developed a Bayesian framework using variational inference to infer Neural ODE parameters at different hierarchical levels. Yin et al. (2021); Kirchmeyer et al. (2022) used hierarchical models, which decompose dynamics into shared and environment-specific components, to enhance forecasting quality and generalization across different parameter regimes of a DS. Similarly, Bird and Williams (2019) introduced a multi-task dynamical systems model where a latent variable encodes task-specific information for sequence generation across different styles/environments. Inubushi and Goto (2020) and Guo et al. (2021) used reservoir computing for transfer learning between chaotic systems, relying on shared reservoir dynamics without explicitly modeling system-specific dynamics. Finally, Desai et al. (2022) proposed a transfer learning approach for PINNs, fine-tuning only the final layer of a pre-trained model for one-shot learning of ODEs, with the physics-informed component embedding physical priors to reduce the amount of required data.
@@
In all these studies, however, only benchmark systems living in similar dynamical regimes (e.g. all exhibiting periodic orbits) were considered, often only assessing forecasting but not generative performance. This is fundamentally different from the settings with complex, chaotic, topologically diverse dynamics and real-world data we consider here. In particular, we explicitly test the case where different underlying system parameters may tap into different dynamical regimes (e.g. periodic vs. non-periodic, chaotic behavior), request our approach to deliver generative DSR models that work for all of those regimes, and represent the underlying systems in a common low-dimensional space spanned by dynamical systems features.
3 Method
3.1 Hierarchization
@@
Assume we have observed multiple, multivariate time series $\bm{x}_{1...T_{\text{max}}^{(j)}}^{(j)}$ of lengths $T_{\text{max}}^{(j)}$, $j=1\dots S$, such as measurements from related physical systems, from multiple subjects in medical studies (Huys et al., 2016; Valton et al., 2020), or repeated measurements from the same subject across trials (Zeidman et al., 2019). While generally the individual multivariate time series may come from any type of system, in the following we will denote these as ‘subjects’ as our main application examples will be human data. Our main goal is to infer subject-specific DSR models, parameterized by $\bm{\theta}^{(j)}_{\text{DSR}}$, i.e. generative models of the latent dynamics underlying subject-specific observations, approximated by a discrete-time recursive (flow) map of the form
$\bm{z}_{t}={\bm{F}}_{\bm{\theta}^{(j)}_{\text{DSR}}}\big{(}\bm{z}_{t-1},\bm{s}_{t}\big{)},$ (1)
@@
where $\bm{s}_{t}$ are possible external inputs (like task stimuli). Observations are related to the dynamical process $\{\bm{z}_{t}\}$ via some parameterized observation function $\bm{x}_{t}=h_{\bm{\theta}_{\text{obs}}^{(j)}}(\bm{z}_{t})$. The differences between subjects are captured within an ideally low-dimensional parameter space, represented here by learnable subject-specific features $\bm{l}^{(j)}\in\mathbb{R}^{N_{\text{feat}}}$ which are mapped onto the parameters $\bm{\theta}_{\text{DSR}}^{(j)}$ of subject-specific DS models through a function parameterized by group-level parameters $\bm{\theta}_{\text{group}}$:
@@
The model Eqs. 1-2 is then trained simultaneously on data from all subjects, where features $\bm{l}^{(j)}$ depend only on time series from subject $j$. The overall approach is illustrated in Fig. 7. In ablation studies, we explored various ways to parameterize the map $G_{\theta_{\text{group}}}$, including flexible and expressive functions like multilayer perceptrons (MLPs) and reparameterizations (see Appx. A.1.3, Table 2), but found that simple linear mappings often gave the best results.
DSR Model
For our experiments we chose a mathematically tractable (cf. Eisenmann et al. (2024)) RNN model introduced for DSR in Hess et al. (2023), the shallow PLRNN (shPLRNN):
$\bm{z}_{t}=\bm{A}\bm{z}_{t-1}+\bm{W}_{1}\phi(\bm{W}_{2}\bm{z}_{t-1}+\bm{h}_{2})+\bm{h}_{1},$ (3)
with latent states $\bm{z}_{t}\in\mathbb{R}^{M}$, diagonal matrix $\bm{A}\in\mathbb{R}^{M\times M}$, connectivity matrices $\bm{W}_{1}\in\mathbb{R}^{M\times L}$ and $\bm{W}_{2}\in\mathbb{R}^{L\times M}$, and thresholds $\bm{h}_{2}\in\mathbb{R}^{L}$ and $\bm{h}_{1}\in\mathbb{R}^{M}$, where $L$ is the dimension of the hidden layer, and rectified-linear-unit nonlinearity $\phi(\cdot)=\textrm{ReLU}(\cdot)=\max(0,\cdot)$.
While our approach is general and can be run with any DSR or TS model, here we spell it out specifically for a hierarchically trained shPLRNN, Eq. 3 (hier-shPLRNN). For this model, the subject-specific DSR model parameters are given by $\bm{\theta}_{\text{DSR}}^{(j)}=\left\{\bm{W}_{1}^{(j)},\bm{W}_{2}^{(j)},\bm{h}_{1}^{(j)},\bm{h}_{2}^{(j)},\bm{A}^{(j)}\right\}$. The subject-level parameters $\bm{\theta}_{\text{DSR}}^{(j)}$ are obtained from the feature vectors $\bm{l}^{(j)}\in\mathbb{R}^{1\times N_{\text{feat}}}$ by linearly projecting group-level parameters into the subject space according to:
@@
$\bm{h}_{1}^{(j)}$ $:=\bm{l}^{(j)}\cdot\bm{P}_{h_{1}},\;\;\bm{h}_{2}^{(j)}:=\bm{l}^{(j)}\cdot\bm{P}_{h_{2}},$
where $\text{mat}(\cdot,m,n)$ denotes the operation of reshaping a vector into an $m\times n$ matrix, and $\bm{P}_{W_{1}}\in\mathbb{R}^{N_{\text{feat}}\times(M\cdot L)}$, $\bm{P}_{W_{2}}\in\mathbb{R}^{N_{\text{feat}}\times(L\cdot M)}$, $\bm{P}_{A}\in\mathbb{R}^{N_{\text{feat}}\times M}$, $\bm{P}_{h_{1}}\in\mathbb{R}^{N_{\text{feat}}\times M}$, and $\bm{P}_{h_{2}}\in\mathbb{R}^{N_{\text{feat}}\times L}$ are group-level parameter matrices learned from all subjects simultaneously. With this, the dynamics of a subject-level shPLRNN, specified through features $\bm{l}^{(j)}$, is given by:
$\bm{z}_{t}=\bm{A}^{(j)}\bm{z}_{t-1}+\bm{W}_{1}^{(j)}\phi(\bm{W}_{2}^{(j)}\bm{z}_{t-1}+\bm{h}_{2}^{(j)})+\bm{h}_{1}^{(j)},$ (4)
3.2 Training Method
Generalized Teacher Forcing (GTF)

### v2

@@ Abstract
In science, we are often interested in obtaining a generative model of the underlying system dynamics from observed time series. While powerful methods for dynamical systems reconstruction (DSR) exist when data come from a single domain, how to best integrate data from multiple dynamical regimes and leverage it for generalization is still an open question. This becomes particularly important when individual time series are short, and group-level information may help to fill in for gaps in single-domain data. Here we introduce a hierarchical framework that enables to harvest group-level (multi-domain) information while retaining all single-domain characteristics, and showcase it on popular DSR benchmarks, as well as on neuroscience and medical data. In addition to faithful reconstruction of all individual dynamical regimes, our unsupervised methodology discovers common low-dimensional feature spaces in which datasets with similar dynamics cluster. The features spanning these spaces were further dynamically highly interpretable, surprisingly in often linear relation to control parameters that govern the dynamics of the underlying system. Finally, we illustrate transfer learning and generalization to new parameter regimes, paving the way toward DSR foundation models.
@@ 1 Introduction
In scientific and medical applications, beyond mere time series forecasting, we are often interested in interpretable and mathematically tractable models that can be used to obtain mechanistic insights into the underlying system dynamics. Such models need to be generative in a dynamical systems sense (Durstewitz et al., 2023), i.e., when simulated, should produce dynamical behavior with the same long-term statistics as the observed system, a property that conventional time series models often lack (Fig. 29, Gilpin (2024)). This is called dynamical systems (DS) reconstruction (DSR), a currently burgeoning field in scientific ML (reviewed in Durstewitz et al. (2023); Gilpin (2024)). Accounting for the fact that different time series may be sampled from fundamentally different dynamical regimes, i.e. with different state space topology and attractor structure, may furthermore endow such models with the ability to transfer knowledge and generalize (Göring et al., 2024). This is particularly important in scientific areas, where data is not only expensive to collect, hence sparse, but also subject to significant variability (Meyer-Lindenberg, 2023; Fechtelpeter et al., 2024). The challenge in these settings is to develop models that can still identify and exploit commonalities across different datasets while being flexible enough to account for individual differences in dynamical regimes.
@@
Hierarchical models provide a natural solution by separating domain-general and domain-specific features. Such models facilitate the transfer of group-level information across individual datasets (Pan and Yang, 2010), avoiding overfitting and improving generalization to unobserved conditions with minimal data (Kirchmeyer et al., 2022). They further explicitly represent inter-domain variability, as often crucial in addressing scientific hypotheses, by unsupervised extraction of a common feature space (Zeidman et al., 2019). While hierarchical modeling is a meanwhile established approach in statistics, including time series analysis (Berliner, 1996; Hyndman et al., 2011), it has so far hardly been applied to the more challenging problem of DSR where we aim for generative models of the system dynamics that can be used for downstream analysis and simulation.
@@
In this work, we propose such a general framework for extracting hierarchical DSR models from multi-domain time series data. Our approach combines low-dimensional, system- or subject-specific feature vectors with high-dimensional, group-level parameters to produce domain-specific DSR models, which can be trained end-to-end using state-of-the-art techniques designed for DSR (Mikhaeil et al., 2022; Hess et al., 2023). We demonstrate the effectiveness of our method in transfer and few-shot learning of DS, where generalization to previously unobserved dynamical regimes becomes possible with minimal new data. More importantly, we find that the low-dimensional subject-specific features learn to represent crucial control parameters of the underlying DS, rendering the subject feature spaces dynamically highly interpretable. We then demonstrate how unsupervised classification in these spaces based on DS features profoundly outperforms other unsupervised methods that harvest more ‘conventional’ time series features, providing a strong point for the DSR perspective on time series analysis.
@@ 2 Related Work
DSR is a rapidly growing field in scientific ML. DSR models may be considered a special class of time series (TS) models that, beyond mere TS prediction, aim to learn surrogate models of the data-generating process from TS observations. A proper DSR model needs to preserve the long-term temporal and geometrical properties of the original DS (Fig. 29), i.e. its vector field topology and attractor structure, which then enables further scientific analysis (Brunton and Kutz, 2019; Durstewitz et al., 2023; Platt et al., 2023; Gilpin, 2024). A variety of DSR methods have been advanced in recent years, either based on predefined function libraries and symbolic regression, such as Sparse Identification of Nonlinear Dynamics (SINDy; Brunton et al. (2016); Kaiser et al. (2018)), based on Koopman operator theory (Azencot et al., 2020; Brunton et al., 2021; Naiman and Azencot, 2021), on universal approximators such as recurrent neural networks (RNNs; Gajamannage et al. (2023); Brenner et al. (2022); Hess et al. (2023)), reservoir computing (RC; Pathak et al. (2017); Platt et al. (2022; 2023)), switching or decomposed linear DS (Linderman et al., 2017; Mudrik et al., 2024a), neural ordinary differential equations (Neural ODEs; Chen et al. (2018); Ko et al. (2023)), or on methods that sit somewhere in between universal and domain-specific, like physics-informed neural networks (PINNs; Raissi et al. (2019)). PINNs, like library-based methods, require sufficient domain knowledge to work well in practice (Fotiadis et al., 2023; Subramanian and Mahadevan, 2023; Mouli et al., 2023). To achieve proper reconstruction of a system’s long-term statistics and attractor geometry, often special, control-theoretic training techniques such as sparse or generalized teacher forcing (Brenner et al., 2022; Mikhaeil et al., 2022; Hess et al., 2023) or particular regularization terms that enforce invariant measures (Platt et al., 2023; Jiang et al., 2023; Schiff et al., 2024) are used. Models using these training techniques currently represent the state of the art in this field.
@@
Hierarchical models for representing multi-level dependencies or nested groups have a long history in statistics and machine learning (Laird and Ware, 1982; Goldstein, 1987; Gelman and Hill, 2006; McCulloch et al., 2011). In the time series domain, Zoeter and Heskes (2003); Bakker and Heskes (2007), for instance, introduced dynamic hierarchical models that combine individual time series with group-level dependencies. Hierarchical models have also been applied for forecasting chaotic time series (Matsumoto et al., 1998; Hyndman et al., 2011; Xu et al., 2020), or for extracting interpretable summary statistics from time series to capture inter-individual differences (Akintayo and Sarkar, 2018; Yingzhen and Mandt, 2018). Much less work exists in the direction of hierarchical DS modeling. LFADS (Pandarinath et al., 2018) and CrEIMBO (Mudrik et al., 2024b) were designed for inferring DS with neuroscience applications in mind, and can integrate observations obtained across different sessions. Roeder et al. (2019), more generally, developed a Bayesian framework using variational inference to infer Neural ODE parameters at different hierarchical levels. Yin et al. (2021); Kirchmeyer et al. (2022) used hierarchical models, which decompose dynamics into shared and environment-specific components, to enhance forecasting quality and generalization across different parameter regimes of a DS. Similarly, Bird and Williams (2019) introduced a multi-task DS model where a latent variable encodes task-specific information for sequence generation across different styles/environments. Inubushi and Goto (2020) and Guo et al. (2021) used reservoir computing for transfer learning between chaotic systems, but without explicitly modeling system-specific dynamics. Finally, Desai et al. (2022) proposed a transfer learning approach for PINNs, fine-tuning only the final layer of a pre-trained model for one-shot learning of ODEs.
@@
In all these studies, however, only benchmark systems living in similar dynamical regimes (e.g. all exhibiting periodic orbits) were considered, often only assessing forecasting but not generative performance. This is fundamentally different from the settings with complex, chaotic, topologically diverse dynamics and real-world data we consider here, with a focus on generative and long-term performance across different dynamical regimes.
3 Methods
3.1 Hierarchical DSR model
@@
Assume we have observed multiple, multivariate time series $\bm{x}_{1...T_{\text{max}}^{(j)}}^{(j)}$ of lengths $T_{\text{max}}^{(j)}$, $j=1\dots S$, such as measurements from related physical systems or from multiple subjects in medical studies. While generally the individual multivariate time series may come from any type of system, in the following we will denote these as ‘subjects’ as our main application examples will be human data. Our main goal is to infer subject-specific DSR models, parameterized by $\bm{\theta}^{(j)}_{\text{DSR}}$, i.e. generative models of the latent dynamics underlying subject-specific observations. We approximate the dynamics by a discrete-time recursive (flow) map of the form
$\bm{z}^{(j)}_{t}={\bm{F}}_{\bm{\theta}^{(j)}_{\text{DSR}}}\big{(}\bm{z}^{(j)}_{t-1},\bm{s}^{(j)}_{t}\big{)},$ (1)
where $\bm{s}_{t}$ are possible external inputs (like task stimuli). Observations are related to the dynamical process $\{\bm{z}_{t}\}$ via some parameterized observation function
$\hat{\bm{x}}_{t}^{(j)}=h_{\bm{\theta}_{\text{obs}}^{(j)}}(\bm{z}^{(j)}_{t}).$ (2)
@@
The differences between subjects are captured within a low-dimensional parameter space, represented here by learnable subject-specific features $\bm{l}^{(j)}\in\mathbb{R}^{N_{\text{feat}}}$. These are mapped onto the parameters $\bm{\theta}_{\text{DSR}}^{(j)}$ of subject-specific DSR models through a function parameterized by group-level parameters $\bm{\theta}_{\text{group}}$ common to all subjects:
@@
The overall approach is illustrated in Fig. 1.
Figure 1: Illustration of the hierarchization framework.
For our experiments we chose for ${\bm{F}}_{\bm{\theta}^{(j)}_{\text{DSR}}}$ (Eq. 1) an RNN model introduced for DSR in Hess et al. (2023), the shallow PLRNN (shPLRNN),
$\bm{z}^{(j)}_{t}={\bm{F}}_{\bm{\theta}^{(j)}_{\text{DSR}}}\big{(}\bm{z}_{t-1}^{(j)}\big{)}=\bm{A}^{(j)}\bm{z}^{(j)}_{t-1}+\bm{W}_{1}^{(j)}\phi(\bm{W}_{2}^{(j)}\bm{z}^{(j)}_{t-1}+\bm{h}_{2}^{(j)})+\bm{h}_{1}^{(j)},$ (4)
with latent states $\bm{z}^{(j)}_{t}\in\mathbb{R}^{M}$, subject-specific parameters $\bm{\theta}_{\text{DSR}}^{(j)}=\{\bm{W}_{1}^{(j)}\in\mathbb{R}^{M\times L},\allowbreak\bm{W}_{2}^{(j)}\in\mathbb{R}^{L\times M},\allowbreak\bm{h}_{1}^{(j)}\in\mathbb{R}^{M},\allowbreak\bm{h}_{2}^{(j)}\in\mathbb{R}^{L},\bm{A}^{(j)}\in\texttt{diag}(\mathbb{R}^{M})\}$, and nonlinearity $\phi(\cdot)=\textrm{ReLU}(\cdot)=\max(0,\cdot)$, which makes the model piecewise-linear and thus mathematically tractable (Eisenmann et al., 2024). We also tested other models for ${\bm{F}}_{\bm{\theta}^{(j)}_{\text{DSR}}}$, including LSTMs (Hochreiter and Schmidhuber, 1997; Vlachas et al., 2018) and vanilla RNNs trained by GTF (Hess et al., 2023), and found that the results reported below for the shPLRNN stayed essentially the same, see Appx. A.1.6 and Table 3.
Likewise, for mapping the latent states $\bm{z}_{t}$ to the observations $\bm{x}_{t}$, Eq. 2, we checked different options, see Appx. A.1.6. Taking simply the identity mapping, $\hat{\bm{x}}_{t}^{(j)}=h_{\bm{\theta}_{\text{obs}}^{(j)}}(\bm{z}_{t}^{(j)})=\bm{z}_{t}^{(j)}$, as in Hess et al. (2023), worked about best, presumably because a too expressive observation model can remove the burden from the DSR model to account for the actual dynamics (Hess et al., 2023).
Finally, in ablation studies we also explored various ways to parameterize the map $G_{\bm{\theta}_{\text{group}}}$ in Eq. 3, including flexible and expressive functions like multilayer perceptrons (MLPs) and reparameterizations (see Appx. A.1.7, Table 4), but found that simple linear mappings often gave the best results. Specifically, for the shPLRNN in Eq. 4, the subject-level parameters $\bm{\theta}_{\text{DSR}}^{(j)}$ are obtained from the feature vectors $\bm{l}^{(j)}\in\mathbb{R}^{1\times N_{\text{feat}}}$ by linearly projecting common group-level parameters  $\bm{\theta}_{\text{group}}=\{\bm{P}_{W_{1}}\in\mathbb{R}^{N_{\text{feat}}\times(M\cdot L)},\bm{P}_{W_{2}}\in\mathbb{R}^{N_{\text{feat}}\times(L\cdot M)},\bm{P}_{h_{1}}\in\mathbb{R}^{N_{\text{feat}}\times M},\bm{P}_{h_{2}}\in\mathbb{R}^{N_{\text{feat}}\times L},\bm{P}_{A}\in\mathbb{R}^{N_{\text{feat}}\times M}\}$ into the subject parameter space according to:
@@
$\bm{h}_{1}^{(j)}$ $:=[\bm{l}^{(j)}\cdot\bm{P}_{h_{1}}]^{T},\;\;\bm{h}_{2}^{(j)}:=[\bm{l}^{(j)}\cdot\bm{P}_{h_{2}}]^{T},$
where $\text{mat}(\cdot,m,n)$ denotes the operation of reshaping a vector into an $m\times n$ matrix.
3.2 Training Procedure
All model parameters, i.e. common group-level parameters $\bm{\theta}_{\text{group}}$ (matrices $\bm{P}$ above) and subject-specific features $\bm{l}^{(1:S)}$, are optimized simultaneously (end-to-end) by minimizing the total loss
$\mathcal{L}=\sum_{j=1}^{S}\sum_{t=1}^{T_{\text{max}}^{(j)}}\frac{1}{2}\log|\bm{\Sigma}^{(j)}|+\frac{1}{2}(\bm{x}_{t}^{(j)}-\hat{\bm{x}}_{t}^{(j)})^{T}(\bm{\Sigma}^{(j)})^{-1}(\bm{x}_{t}^{(j)}-\hat{\bm{x}}_{t}^{(j)}).$ (5)
The total loss thus contains additional parameters $\bm{\Sigma}^{(j)}$ trained jointly with $\bm{\theta}_{\text{group}}$ and $\bm{l}^{(1:S)}$, which are subject-specific diagonal scaling matrices introduced to deal with data that is on different scales across subjects (e.g. for observations of the Lorenz-63 system in the cyclic vs. chaotic regime). For gradient propagation, note that predictions $\hat{\bm{x}}_{t}^{(j)}$ in Eq. 5 are linked via Eq. 2 to the latent model, Eq. 4, whose parameters $\bm{\theta}_{\text{DSR}}^{(j)}$ are in turn obtained from $\bm{\theta}_{\text{group}}$ and $\bm{l}^{(j)}$ via Eq. 3, see Fig. 8. For the group-level matrices $\bm{P}_{\cdot}\in\bm{\theta}_{\text{group}}$, Xavier uniform initialization (Glorot and Bengio, 2010) was used and an $L_{2}$ norm added to the loss, see Appx. A.1 for more details.